# Extensions of $\mathcal{KL}$ and Lyapunov Functions for Discrete-time Dynamical System Peaks Analysis

### v3

@@ 6 A Technical Summary
As we have $f(x)>x$ for all $x>0$, we have $f(1)>1$ and thus the interval $[1,f(1))$ is nonempty. We define $h:[1,f(1))$ continuous, strictly increasing whose values are strictly positive such that $h(1)=1$ and $\lim_{x\mapsto f(1)^{-}}h(x)=\gamma$. We could take for example $h$ the affine map linking $(1,1)$ to $(f(1),\gamma)$. Now, as the sequence $(f^{k}(1))_{k}$ is strictly increasing and unbounded (if $M=\sup_{k\in\mathbb{N}}f^{k}(1)>0$ is finite then as $f$ is increasing and continuous $f(M)=M$ but $M<f(M)$ by assumption on $f$). As $f$ is invertible on $\mathbb{R}_{+}$ and we have $f^{-1}(x)<x$ for all $x>0$, the sequence $(f^{-k}(1))_{k}$ is strictly decreasing and positive and converges to 0 as $k$ tends to +$\infty$+ ∞. Hence $\mathbb{R}_{+}^{*}$ is union of disjoint intervals $[f^{k}(1),f^{k+1}(1))$ where $k\in\mathbb{Z}$. Then for all $x>0$, there exists a unique $k\in\mathbb{Z}$ such that $x\in[f^{k}(1),f^{k+1}(1))$. We define a function $g$ on $\mathbb{R}_{+}$ as follows $g(0)=0$ and for all $x>0$, $g(x):=\gamma^{k}h(f^{-k}(x))$ where $k$ is the unique integer such that $x\in[f^{k}(1),f^{k+1}(1))$. As $g$ is equal to $h$ on $[1,f(1)]$, then $g$ is strictly increasing, continuous and strictly positive on $[1,f(1)]$. Then, on the open interval $(f^{k}(1),f^{k+1}(1))$ for $k\neq 0$, $g$ is the composition $t\mapsto\gamma^{k}t$, $h$ and $f^{-k}$ and thus is strictly increasing, continuous and strictly positive on $(f^{k}(1),f^{k+1}(1))$ for all $k\in\mathbb{Z}$. Now, for all $k\neq 0$, $\lim_{y\to{f^{k}(1)}^{-}}g(y)=\lim_{y\to{f^{k}(1)}^{-}}\gamma^{k-1}h(f^{-k+1}(y))=\lim_{z=f^{-k}(y)\to 1^{-}}\gamma^{k-1}h(f(z))=\gamma^{k-1+1}=\gamma^{k}$. We also have $\lim_{y\to{f^{k}(1)}^{+}}g(y)=\lim_{y\to{f^{k}(1)}^{+}}\gamma^{k}h(f^{-k}(y))=\lim_{z\to 1^{-}}\gamma^{k}h(z)=\gamma^{k}h(1)=\gamma^{k}$. We conclude that $g$ is continuous on $\mathbb{R}_{+}^{*}$. Now, we prove that $g$ is continuous at 0 (at right). Let us consider a sequence $(x_{n})_{n\in\mathbb{N}}$ of positive reals which converges to 0. Let $n\in\mathbb{N}$. There exists a unique $k_{n}\in\mathbb{Z}$ such that $f^{k_{n}}(1)\leq x_{n}\leq f^{k_{n}+1}(1)$. We get $1\leq f^{-k_{n}}(x_{n})\leq f(1)$ and as $h$ is increasing on $[1,f(1))$ and continuous (at left) at $f(1)$, $1=h(1)\leq h(f^{-k_{n}}(x_{n}))\leq h(f(1))=\gamma$. Finally, $\gamma^{k_{n}}\leq g(x_{n})=\gamma^{k_{n}}h(f^{-k_{n}}(x_{n}))\leq\gamma^{k_{n}+1}$. As $(x_{n})_{n\in\mathbb{N}}$ tends to 0, the sequence $(k_{n})_{n\in\mathbb{N}}$ tends to $\infty$- ∞ and we conclude that $(g(x_{n}))_{n\in\mathbb{N}}$ tends to $0=g(0)$ and $g$ is continuous at 0. We have already proved that $g$ is strictly increasing on each interval $[f^{k}(1),f^{k+1}(1))$. It is easy to see that it is also the case when we pick two reals $x,y$ in two different intervals : $x<y$ implies $g(x)<g(y)$.
@@
Finally, we have to prove that $g(f(x))=\gamma g(x)$ for all $x\in\mathbb{R}_{+}$. As $f(0)=0$, this holds for $x=0$. Now let us take $x>0$. This real $x$ belongs to some interval $[f^{k}(1),f^{k+1}(1))$. Thus, $f(x)\in[f^{k+1}(1),f^{k+2}(1))$ and $g(f(x))=\gamma^{k+1}h(f^{-k-1}(f(x))=\gamma\gamma^{k}h(f^{-k}(x))=\gamma g(x)$. This ends the proof. ∎
Proof of Lemma 5
Lemma 5.

### v4

@@ A.1 Additional proofs for Opt-Lyapunov definition
As we have $f(x)>x$ for all $x>0$, we have $f(1)>1$ and thus the interval $[1,f(1))$ is nonempty. We define $h:[1,f(1))$ as continuous and strictly increasing, with strictly positive values such that $h(1)=1$ and $\lim_{x\mapsto f(1)^{-}}h(x)=\gamma$. For example, we could take $h$ to be the affine map linking $(1,1)$ to $(f(1),\gamma)$. Now, as the sequence $(f^{k}(1))_{k}$ is strictly increasing and unbounded (if $M=\sup_{k\in\mathbb{N}}f^{k}(1)>0$ is finite, then as $f$ is increasing and continuous, $f(M)=M$ but $M<f(M)$ by the assumption on $f$). As $f$ is invertible on $\mathbb{R}_{+}$ and we have $f^{-1}(x)<x$ for all $x>0$, the sequence $(f^{-k}(1))_{k}$ is strictly decreasing and positive and converges to 0 as $k$ tends to $\infty$. Hence $\mathbb{R}_{+}^{*}$ is union of disjoint intervals $f^{k}(1),f^{k+1}(1))$ where $k\in\mathbb{Z}$. Then for all $x>0$, there exists a unique $k\in\mathbb{Z}$ such that $x\in[f^{k}(1),f^{k+1}(1))$. We define a function $g$ on $\mathbb{R}_{+}$ as follows $g(0)=0$ and for all $x>0$, $g(x):=\gamma^{k}h(f^{-k}(x))$ where $k$ is the unique integer such that $x\in[f^{k}(1),f^{k+1}(1))$. As $g$ is equal to $h$ on $[1,f(1)]$, then $g$ is strictly increasing, continuous, and strictly positive on $[1,f(1)]$. Then, on the open interval $(f^{k}(1),f^{k+1}(1))$ for $k\neq 0$, $g$ is the composition $t\mapsto\gamma^{k}t$, $h$ and $f^{-k}$ and thus is strictly increasing, continuous and strictly positive on $(f^{k}(1),f^{k+1}(1))$ for all $k\in\mathbb{Z}$. Now, for all $k\neq 0$, $\lim_{y\to{f^{k}(1)}^{-}}g(y)=\lim_{y\to{f^{k}(1)}^{-}}\gamma^{k-1}h(f^{-k+1}(y))=\lim_{z=f^{-k}(y)\to 1^{-}}\gamma^{k-1}h(f(z))=\gamma^{k-1+1}=\gamma^{k}$. We also have $\lim_{y\to{f^{k}(1)}^{+}}g(y)=\lim_{y\to{f^{k}(1)}^{+}}\gamma^{k}h(f^{-k}(y))=\lim_{z\to 1^{-}}\gamma^{k}h(z)=\gamma^{k}h(1)=\gamma^{k}$. We conclude that $g$ is continuous on $\mathbb{R}_{+}^{*}$. Now, we prove that $g$ is continuous at 0 (at right). Let us consider a sequence $(x_{n})_{n\in\mathbb{N}}$ of positive reals that converges to 0. Let $n\in\mathbb{N}$. There exists a unique $k_{n}\in\mathbb{Z}$ such that $f^{k_{n}}(1)\leq x_{n}\leq f^{k_{n}+1}(1)$. We get $1\leq f^{-k_{n}}(x_{n})\leq f(1)$ and as $h$ is increasing on $[1,f(1))$ and continuous (at left) at $f(1)$, $1=h(1)\leq h(f^{-k_{n}}(x_{n}))\leq h(f(1))=\gamma$. Finally, $\gamma^{k_{n}}\leq g(x_{n})=\gamma^{k_{n}}h(f^{-k_{n}}(x_{n}))\leq\gamma^{k_{n}+1}$. As $(x_{n})_{n\in\mathbb{N}}$ tends to 0, the sequence $(k_{n})_{n\in\mathbb{N}}$ tends to -$\infty$ and we conclude that $(g(x_{n}))_{n\in\mathbb{N}}$ tends to $0=g(0)$ and $g$ is continuous at 0. We have already proven that $g$ is strictly increasing on each interval $f^{k}(1),f^{k+1}(1))$. It is easy to see that this is also the case when we pick two reals $x,y$ in two different intervals: $x<y$ implies $g(x)<g(y)$.
@@
Finally, we must prove that $g(f(x))=\gamma g(x)$ for all $x\in\mathbb{R}_{+}$. As $f(0)=0$, this holds for $x=0$. Now, let us take $x>0$. This real $x$ belongs to some interval $f^{k}(1),f^{k+1}(1))$. Thus, $f(x)\in[f^{k+1}(1),f^{k+2}(1))$ and $g(f(x))=\gamma^{k+1}h(f^{-k-1}(f(x))=\gamma\gamma^{k}h(f^{-k}(x))=\gamma g(x)$. This ends the proof. ∎
Lemma A.4.